OpenStudy (anonymous):

An artisan is creating a circular street mural for an art festival. The mural is going to be 50 feet wide. One sector of the mural spans 38 degrees. What is the area of the sector to the nearest square foot?

OpenStudy (anonymous):

Remember formula for area of sector? \[A = \frac{ x }{ 360 } \times \pi r^2\]

OpenStudy (anonymous):

would 50 or 38 go in place of x?

OpenStudy (anonymous):

@whatdoesthismean